What is the change in a) consumption and b) change in savings, if with a marginal propensity to consume (MPC) of 73%, real disposable income (Yd) increases from $215 to $295? Answer:
Answer #1

This is given that a marginal propensity to consume (MPC) of 73%,

MPC=0.73

real disposable income (Yd) increases from $215 to $295

When Yd=215

C=MPC*Yd

=0.73*215

=156.55

Saving=Yd-C

=215-156.55

=58.05

When Yd=295

C=MPC*Yd

=0.73*295

=215.35

Saving=Yd-C

=295-215.35

=79.65

Change in consumption=215.35-156.55

=58.8

Change in saving=79.65-58.05

=21.60